OpenStudy (anonymous):

simplify (5)^-5(5)^7

OpenStudy (anonymous):

is it 25?

OpenStudy (anonymous):

use the rule \[\Large x^p * x^q = x^{(p+q)}\]

OpenStudy (anonymous):

oh so itd be\[5^-5+7=25\]

OpenStudy (anonymous):

you mean 5^2 which makes it 25?

OpenStudy (anonymous):

as 5-5+7 is not 25, but i assume you meant 5^(-5+7) = 5^2 = 25

OpenStudy (anonymous):

which is correct

OpenStudy (anonymous):

\[5^{(-5+7)}=5^2=25\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh i get it now thank you!